Ekmark Corporation uses the following activity rates from its activity-based costing to assign overhead costs to products: Assembling products.................. $6.56 per assembling hour Processing customer orders......... $65.38 per customer order Setting up batches...................... $82.84 per batch Data for one of the company's products follow: Product P59G Number of Assembly Hours.............240 Number of Customer Orders...........48 Number of Batches.......................64 How much overhead cost would be allocated to Product P59G in the 2nd stage using activity based costing system? a. $54,482.56 b. $5,301.76 c. $10,014.40 d. $154.78
Answer #1

C. $10,014.40

Assembling cost = $6.56*240 hours = $$1574.40

Processing cost = $65.38*48 hours = $3138.24

Setting up cost = $82.84*64 = $5301.76

Total product cost = $10014.40